OpenStudy (anonymous):

If f(x) = 2x − 2 and g(x) = 3x − 1, find (f − g)(4).

OpenStudy (lgbasallote):

(f-g)(4) means f(4) - g(4) so evaluate f(4) f(4) = 2(4) - 2 then evaluate g(4) g(4) = 3(4) - 1

OpenStudy (lgbasallote):

tell me what you get for those later